Suppose tan(theta) = 4/3 and csc<0, find tan(theta/2)
1
Expert's answer
2020-11-11T12:49:42-0500

Using half-angle formula.

"\tan\theta=\frac{2t}{1-t^2}" , where "t=\tan\frac{\theta}{2}" .

"\frac{4}{3}=\frac{2t}{1-t^2}\\
4(1-t^2)=6t\\
2(1-t^2)=3t\\
2-2t^2=3t\\
2t^2+3t-2=0\\
2t^2+4t-t-2=0\\
2t(t+2)-1(t+2)=0\\
(2t-1)(t+2)=0\\
2t-1=0 ,t+2=0\\
t=\frac{1}{2}, t=-2"

So,

"\tan\frac{\theta}{2}=\frac{1}{2} or-2"


But,. there is a condition that "\csc\theta<0" .


If "\tan\frac{\theta}{2}=\frac{1}{2}" , then "\csc\theta=\frac{5}{4}>0" . Therefore, this is not the answer.


If "\tan\frac{\theta}{2}=-2," then "\csc\theta=-\frac{5}{4}<0" . Therefore this is the answer.


So, "\tan\frac{\theta}{2}=-2."